OpenStudy (anonymous):

So I need the formula to put in my calculator when finding how many percent $300 is in $60,000

OpenStudy (amistre64):

spose we have a total of T, 1% is simply a division by 100 T/100 = 1% by mutiplying both sides by some value, n, we can determine the n% of T nT/100 = n% so the question is, when does nT/100 = 300, given that T=60000

OpenStudy (anonymous):

But how would I type this in my calculator, is it 300/100*60000?

OpenStudy (amistre64):

let do some algebra :) nT/100 = 300 nT = 300*100 n = 300*100/T

OpenStudy (amistre64):

soo n = 30000 ------ 60000 n = 3/6 = 1/2. 300 is 1/2% of 60000

OpenStudy (anonymous):

Oh okayy so will the answer be 0.5?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

Thanksss !!

OpenStudy (amistre64):

youre welcome